OpenStudy (karolynxx):

Which equation shows the point-slope form of the line that passes through (3, 2) and has a slope of 1/3? a. y + 2 =1/3(x + 3) b. y – 2 = 1/3(x – 3) c. y + 3 =1/3 (x + 2) d. y – 3 = 1/3(x – 2)

OpenStudy (anonymous):

What are the options?

OpenStudy (karolynxx):

i just edited the options

OpenStudy (alexandervonhumboldt2):

point slope form is y-y1=m*(x-x1) where (x1, y1) is a coordinate and m is slope

OpenStudy (anonymous):

The way to solve this is plugging in the data on the original equation of a linear function and then convert it to the point slop form the original equation is y=mx+n where m is the slope and n the y-intercept the point slope form is y-y1=m*(x-x1)

OpenStudy (karolynxx):

so its b or c?

OpenStudy (anonymous):

option B im going with

OpenStudy (anonymous):

yhea its b